OpenStudy (anonymous):

The population of bacteria in a Petri dish increases by a a factor of 10 every hour. If there are initially 20 bacteria in the dish, how long will it take before the population increases to 60 000 to the nearest tenth of an hour?

OpenStudy (anonymous):

\[20\times 10^x=60,000\]

OpenStudy (anonymous):

\[10^x=3000\]

OpenStudy (anonymous):

\[x=?\]

OpenStudy (anonymous):

Would i have to use the log or ln functions since that is what we're learning?

OpenStudy (math&ing001):

Yes, you'll need to use this : \[\log(10^{x})=x\]

OpenStudy (anonymous):

oh alright, thank you :)